cassidy is 3 times as old as drew. in 8 years cassidy will only be twice as old as drew. how old is cassidy right now?

Answer by oberobic(2304)

c= cassidy's current age
c+8 = cassidy's age in 8 yr
.
d = drew's current age
d+8 = drew's age in 8 yr
.
c = 3*d
.
c+8 = 2*(d+8)
.
substitute c = 3d
.
3d + 8 = 2d + 16
.
d = 8
.
c = 3(8) = 24
.
Check the relationship in 8 years.
.
c+8 = 32
d+8 = 16
So c = 2d, which is correct.
.
Answer: Cassidy is 24; Drew is 8.
.
Done.